# Hadamard Representations: Augmenting Hyperbolic Tangents in RL

### v3

@@ Abstract
Activation functions are one of the key components of a deep neural network. The most commonly used activation functions can be classed into the category of continuously differentiable (e.g. tanh) and piece-wise linear functions (e.g. ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning (e.g. measured by the number of dead neurons and the effective rank). In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to piece-wise linear functions. We provide insights into the vanishing gradients associated with the former, and show that the dying neuron problem is not exclusive to ReLU’s. To alleviate vanishing gradients and the resulting dying neuron problem occurring with continuously differentiable activations, we propose a Hadamard representation. Using deep Q-networks, proximal policy optimization and parallelized Q-networks in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
@@ 1 Introduction
Figure 1: Normalized performance training DQN in the Atari domain where the nonlinear activation function of the representation, defined as the final hidden layer, is either ReLU or tanh. A strong performance discrepancy due to a relatively small architectural change can be observed.
@@
Out of all activation functions, the Rectified Linear Unit (ReLU) (Fukushima, 1969; Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016). The strength of the ReLU activation lies in its ability to naturally avoid vanishing gradients when used in deeper networks, in contrast to the continuously differentiable activation functions, such as the sigmoid and the hyperbolic tangent (Glorot & Bengio, 2010).
@@
A common drawback of using the ReLU activation is its limited expressivity in the context of shallow networks (see Fig. 2), as well as the phenomenon known as the dying ReLU problem (He et al., 2015; Lu et al., 2019). As training progresses, the number of dying ReLU’s tend to increase, resulting in a dying network and loss of network capacity (Dubey et al., 2022).
@@
In reinforcement learning (RL) (Sutton & Barto, 2018), the dying neuron phenomenon is more prevalent than in supervised learning due to the use of non-stationary targets (Sokar et al., 2023). However, even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018). Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not commonly used in RL (see Fig. 1). However, one might argue that their symmetrical, bounded shape and smooth gradient landscape offer optimization advantages that the ReLU lacks. Recent findings also indicate that a hidden layer activated by a hyperbolic tangent displays a high effective rank and thus a high layer expressivity (Kumar et al., 2021; Gulcehre et al., 2022). Despite being a theoretically sound candidate, its lack of success in RL has not been thoroughly investigated.
@@
We therefore aim to provide insights into the hyperbolic tangent’s suboptimality, revealing that the vanishing gradient problem leads to dying neurons and under-utilization of the full network capacity. Based on these insights, we mitigate said effects by augmenting the original architecture. Specifically, we focus our research on the activations of an encoder, defined as the hidden layer, and provide an alternative to the conventional parameterization of this representation. Our contributions can be summarized as follows:
@@
  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem, but have a more profound effect on performance.
@@
  • A Hadamard representation (HR) is proposed, defining a latent representation as the Hadamard product of two separate, individually parameterized activation vectors.
@@
  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, using Hadamard representations yields notable performance gains in multiple algorithms in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
@@ 2 Preliminaries
We consider an agent acting within its environment, where the environment is modeled as a discrete Markov Decision Process (MDP) defined as a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$. $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the environment’s transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{R}$ is the environment’s reward function and $\gamma\in[0,1)$ is the discount factor. A replay buffer $B$ is used to store visited states $s_{t}\in\mathcal{S}$ that were followed by actions $a_{t}\in\mathcal{A}$ and resulted in the rewards $r_{t}\in\mathcal{R}$ and the next states $s_{t+1}$. One entry in $B$ contains a tuple of past experience $(s_{t},a_{t},r_{t},s_{t+1})$. The agent’s goal is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expectation of the discounted return $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\tau}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{t}=s]$, where $\tau$ is a trajectory following the policy $\pi$.
Furthermore, we examine the setting where a high-dimensional state ($s_{t}\in\mathbb{R}^{v}$) is compressed into lower-dimensional activations $z_{t}\in\mathcal{Z}=\mathbb{R}^{w}$ where we call $\mathcal{Z}$ the representation space with $w\leq v$. This is done by means of a neural network encoding $e:\mathcal{S}\rightarrow\mathcal{Z}$ where $e$ represents the encoder.
@@ 3 Augmenting Hyperbolic Tangents
Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior. When using any activation function in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, is saturated or dying if:
$\alpha_{i}\approx\Omega\;\forall\;s_{t}\in B$ (1)
@@
Where $\Omega$ represents the saturation value and $s_{t}$ is an observation in buffer $B$. In practice, a mini-batch of observations is evaluated instead of the whole dataset in $B$. For the hyperbolic tangent, given that it is an asymptotic function near its saturation point, an approximate equality must be considered, as the classification of its saturation will always remain qualitative ($|\alpha_{i}|\neq 1\;\forall\;s_{t}\in B$) . To approximate the condition given in Eq. 1, the amount of dying hyperbolic tangents is calculated by using a kernel density estimation (KDE) (Silverman, 1986) on the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the activation layer. In order to visualize activations in a hidden layer, a fixed subset of the KDE’s of the neurons $\alpha_{i}$ is taken. A clear visualization of dying hyperbolic tangents during training in the Atari Breakout environment can be seen by analyzing sixteen individual neuron KDE’s in Fig. 3.
(a) Tanh - $10^{6}$ iterations
(b) Tanh - $10^{7}$ iterations
Figure 3: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN Mnih et al. (2015) in the Breakout environment using a hyperbolic tangent activation for $z_{t}$. Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$. Red outlines represent dying neurons, where a near infinite sized density spike occurs at either 1 or -1.
More details on the KDE calculation and dying neuron classification can be found in Appendix B.
@@
As Fig. 3 indicates that the activation of $z_{t}$ with a hyperbolic tangent leads to saturation and dying neurons, an augmentation of the representation architecture is proposed. In the conventional encoder setting, the networks’ final hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{i}$ and $B_{i}$ representing weight and bias parameters, the function $f()$ representing a nonlinear activation function while $x$ is the set of activations from the previous layer. In order to reduce the information dependence on a single set of neurons, we propose using a Hadamard representation that augments the original representation with a parallel representation layer $z^{*}$. This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015), or as an augmented version of the Gated Linear Unit (GLU) (Dauphin et al., 2017). The final representation is defined as the Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$. A visualization of the proposed architecture can be found in Fig. 4.
Figure 4: A visualisation of the Hadamard Representation architecture combined with the nature DQN architecture applied on a snapshot of the ‘SpaceInvaders’ Atari environment (Mnih et al., 2015). Vertical and horizontal bars represent convolutional and linear layers, respectively. A parallel hidden layer providing independently parameterized activations is integrated, where the Hadamard product of the two hidden layers represents the final latent representation $z_{t}$ used for downstream learning.
@@
Our key hypothesis is that the Hadamard representation can prevent saturation, hence alleviating vanishing gradients and dying neurons (See Eq. 1). To support our hypothesis, we investigate the derivative of a product of two functions. For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}g^{\prime}(x)h(x)}}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}g(x)h^{\prime}(x)}}$. In the context of using a sigmoid activation function for $f(x)$, the derivative of $z(x)$ becomes:
$z^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})}}\color[rgb]{0,0,0}$
$+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))}}$
@@
$z^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})}}\color[rgb]{0,0,0}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\text{sech}^{2}(A_{2}x+B_{2})\tanh(A_{1}x+B_{1})}}$
In this context, $\operatorname{sech}^{2}$ is the derivative of the hyperbolic tangent function. Unlike the sigmoid, the hyperbolic tangent saturates to nonzero values, ensuring that if and only if both parts are saturated, product saturation occours. Thus, when $g(x)$ saturates, $h(x)$ still keeps a non-trivial gradient in the product, providing a mechanism to avoid vanishing gradients. We visualize the kernel densities during training with a Hadamard representation in Fig. 5. The individual representations before taking the Hadamard product can be found in Appendix B.

### v4

@@ Abstract
Activation functions are one of the key components of a deep neural network. The most commonly used activation functions can be classed into the category of continuously differentiable (e.g. tanh) and piece-wise linear functions (e.g. ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning. In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to piece-wise linear functions. We show that the dying neuron problem in RL is not exclusive to ReLUs and actually leads to additional problems in the case of continuously differentiable activations such as tanh. To alleviate the dying neuron problem with these activations, we propose a Hadamard representation that unlocks the advantages of continuously differentiable activations. Using DQN, PPO and PQN in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Median Human-Normalized performance training PQN in the Atari domain where the activation function of the hidden layers is changed. In PQN, all activations are already layer-normalized. A massive performance discrepancy in performance can be observed when selecting different activation functions. Notably, in Atari, the application of a Hadamard representation with hyperbolic tangent leads to over 100% performance gains. The Hadamard representation is not suitable for the ReLU activation, as it amplifies dying neurons by taking the product of sparse activations.
@@
The Rectified Linear Unit (ReLU) (Fukushima, 1969; Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016). The strength of the ReLU activation lies in its ability to naturally avoid vanishing gradients when used in deeper networks, in contrast to the continuously differentiable activation functions, such as the sigmoid and the hyperbolic tangent (Glorot & Bengio, 2010).
@@
A common drawback of using the ReLU activation is its limited expressivity in the context of shallow networks (see Fig. 2), as well as the phenomenon known as the dying ReLU problem (He et al., 2015; Lu et al., 2019). As training progresses, the number of dying ReLUs tend to increase, resulting in a dying network and loss of network capacity (Dubey et al., 2022).
@@
In reinforcement learning (RL) (Sutton & Barto, 2018), the dying neuron phenomenon is much more prevalent than in supervised learning due to the use of non-stationary targets (Sokar et al., 2023). However, even though training results in a large number of dying ReLUs (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018). Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not favored in RL (see Fig. 1). However, one might argue that their symmetrical, bounded shape and smooth gradient landscape offer optimization advantages that the ReLU lacks. Recent findings also indicate that a hidden layer activated by a hyperbolic tangent displays a high effective rank and thus a high layer expressivity (Kumar et al., 2021; Gulcehre et al., 2022). Despite being a theoretically sound candidate, its lack of success in RL has not been thoroughly investigated.
@@
This paper provides insights into the hyperbolic tangent’s suboptimality, revealing that RL gradients lead to bias-inducing dying neurons and under-utilization of the full network capacity. Based on these insights, we mitigate said effects by augmenting the original hidden layer architecture. Specifically, we provide an alternative to the conventional parameterization of these layers. Our contributions can be summarized as follows:
@@
  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem, and argue how they inherently have a more profound effect on performance.
@@
  • A Hadamard representation (HR) is proposed, defining a hidden layer as the Hadamard product of two separate, individually parameterized activation vectors.
@@
  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, Hadamard representations yields notable performance gains in multiple algorithms in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
2 Related Work
This section provides related work from the perspective of: (i) the loss of capacity through learning neural networks and (ii) the effect of different network architectures in RL.
Network Capacity in RL.
Liu et al. (2019) investigated the need for sparse representations in the continuous control domain. Gulcehre et al. (2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the representation, and found that hyperbolic tangent representations generally maintain high rank while not suffering strongly from rank decay as training continues. Related work used normalization techniques and action penalization to counteract high variance in pixel-based robotic control (Bjorck et al., 2022). Other work by Lyle et al. (2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays. Further research by Nikishin et al. (2022) used network resets to counteract the primacy bias and (Sokar et al., 2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting. Nikishin et al. (2023) further studied plasticity injection for long-term training and Delfosse et al. (2024) applied rational activations (Molina et al., 2019) in RL to increase plasticity. Concurrent work by Dohare et al. (2024) used continual back-propagation to further alleviate plasticity loss. In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024). This provides insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
Network Architecture in RL
The origin of network optimization problems with hyperbolic tangents and sigmoids were empirically investigated by Glorot & Bengio (2010), where, according to the authors, a lot of mystery still surrounds the subject. Work by (Srivastava et al., 2015) in supervised learning first looked at the idea of using products of hidden layers together with a ’gate’ that determined the amount of information flow (Hochreiter & Schmidhuber, 1997). Using these ideas, the Resnet was invented (He et al., 2016) and also showed strong performance in combination with RL (Espeholt et al., 2018). Further work by Henderson et al. (2018) showed differences in RL performances over different network architectures and nonlinear activations. Work by Abbas et al. (2023) successfully applied ReLU concatenation (Shang et al., 2016) to improve continual learning while keeping a similar performance when training from scratch. Finally, recent work by Grooten et al. (2024) investigated raw pixel masking for distractions in RL using a parallel CNN input layer.
@@ 3 Preliminaries
(a) Tanh - $10^{6}$ iterations
(b) Tanh - $10^{7}$ iterations
Figure 3: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN (Mnih et al., 2015) in the Breakout environment using a hyperbolic tangent activation for $z_{t}$. Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$. Red outlines represent dying neurons, where a near infinite sized density spike occurs at either 1 or -1.
We consider an agent acting within its environment as a discrete Markov Decision Process (MDP) defined as a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$. $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{P}(\mathcal{S})$ is the environment’s transition function, $R:\mathcal{S}~{}\times~{}\mathcal{A}~{}\rightarrow~{}\mathcal{R}$ is the environment’s reward function and $\gamma\in[0,1)$ is the discount factor. A replay buffer $B$ is used to store visited states $s_{t}\in\mathcal{S}$ that were followed by actions $a_{t}\in\mathcal{A}$ and resulted in the rewards $r_{t}\in\mathcal{R}$ and the next states $s_{t+1}$. One entry in $B$ contains a tuple of past experience $(s_{t},a_{t},r_{t},s_{t+1})$. The agent’s goal is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expectation of the discounted return $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\tau}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{t}=s]$, where $\tau$ is a trajectory following the policy $\pi$.
@@ 4 Augmenting Hyperbolic Tangents
Figure 4: 10M iterations (40M frames) training DQN with a hyperbolic tangent activation in the Seaquest environment. The average contribution to the Q-values of the live and dead neurons in the final hidden layer is observed. If a neuron dies, it retains the same value for any input observation, but a multiplication of the nonzero saturation value with its outgoing weights implements a substantial ‘hidden’ bias on the Q-values.
@@
Although the literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior, particulary in the RL context with the moving target from the Bellman iterations. When using any activation function in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, is saturated or dying if:
$\alpha_{i}\approx\Omega,\;\forall s_{t}\in B$ (1)
Figure 5: A visualisation of the Hadamard representation. Horizontal bars represent weight vectors and $\boldsymbol{z}_{t}$ represents a hidden layer. Between each hidden layer, two parallel independently parameterized activation layers are formed, where the Hadamard product of the two activation layers represents the actual propagated hidden layer.
@@
Where $\Omega$ represents the saturation value and $s_{t}$ is an observation in buffer $B$. In practice, a mini-batch of observations is evaluated instead of the whole dataset in $B$. For the hyperbolic tangent, given that it is an asymptotic function near its saturation point, an approximate equality is considered ($|\alpha_{i}|\neq 1,\;\forall s_{t}\in B$) . To approximate the condition given in Eq. 1, the amount of dying hyperbolic tangents is calculated by using a kernel density estimation (KDE) (Silverman, 1986) on the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the activation layer. In order to visualize activations in a hidden layer, a fixed subset of the KDE’s of the neurons $\alpha_{i}$ is taken. A clear visualization of dying hyperbolic tangents during training in the Atari Breakout environment can be seen by analyzing sixteen individual neuron KDE’s in Fig. 3. Massive KDE spikes at either 1 or -1 represent the absence of neuron variance over a batch of observations, meaning that the neuron has lost its value to propagate useful information. More details on the KDE calculation and dying neuron classification can be found in Appendix B.
Hyperbolic Tangents turn Weights into Biases
Theorem 4.1.
When any set of neurons $\alpha^{j}$ in a hidden layer $z^{j}$ collapses into nonzero values, the output to the next layer effectively changes from ($A^{j}z^{j}+B^{j}$) to ($A_{-*}^{j}z_{i-*}^{j}+B^{j+1}$ + $A_{*}^{j}z_{i*}^{j}$), where $A_{-*}^{j}z_{-*}^{j}$ represent the active neurons multiplied by their corresponding forward-connected weights and $A_{*}^{j}z_{i*}^{j}~{}=~{}B_{*}^{j+1}$ represent the dead neurons multiplied by their corresponding weights, resulting in the hidden bias $B_{*}^{j+1}$.
Proof.
Let us consider a set of neurons $\alpha^{j}_{i}$ and forward connected weights $\boldsymbol{w}^{j}_{\alpha_{i}}$ in the hidden layer $z^{j}$. The influence of these neurons on the next hidden layer $z^{j+1}$ is calculated as:
$z^{j+1}=\sum_{i}\alpha^{j}_{i}\boldsymbol{w}^{j}_{\alpha_{i}}+B^{j+1}.$ (2)
If the set of neurons dies and collapses into 0 ($\alpha^{j}_{i}=0,~{}\forall s\in\mathcal{S}$), which occurs when using ReLU activations, the influence on the next layer becomes 0, representing basic pruning:
$\sum_{i}0\cdot\boldsymbol{\bar{w}}^{j}_{\alpha_{i}}=0,\quad\forall s\in\mathcal{S}.$ (3)
Where $\boldsymbol{\bar{w}}^{j}_{\alpha_{i}}$ is the set of weights connected to the dying neurons. However, for a hyperbolic tangent activation, if the set of neurons saturate into either -1 or 1 ($\alpha^{j}_{i}=\{1,-1\},\forall s\in\mathcal{S}$), the output is
$\sum_{i}\{1,-1\}\cdot\boldsymbol{\bar{w}}^{j}_{\alpha_{i}}=B_{*}^{j+1},\quad\forall s\in\mathcal{S}.$ (4)
As a result, the weight vector $\boldsymbol{\bar{w}}^{j}_{\alpha_{i}}$ corresponding to the dying activations only influences a bias $B^{j+1}_{*}$ on the next layer. Note that the bias $B^{j+1}_{*}$ is constant for any input observation. ∎
This emergent bias can hinder the optimization process by introducing unintended fixed contributions to a networks’ hidden layers or output layer, reducing the flexibility of the network’s representations and potentially reducing its performance in RL or supervised learning. For example, in Atari’s Seaquest environment, we can see that this bias is substantial and nonzero (see Fig. 4).
@@
As Fig. 3 indicates that the activation of $z_{t}$ with a hyperbolic tangent leads to saturation and dying neurons, an augmentation of the representation architecture is proposed. In the conventional encoder setting, a network’s hidden layer can be defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{i}$ and $B_{i}$ representing weight and bias parameters, the function $f()$ representing a nonlinear activation function while $x$ is the set of activations from the previous layer. In order to reduce the information dependence on a single set of neurons, we propose using a Hadamard representation that augments the original representation with a parallel representation layer $z^{*}$. This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015), or as an augmented version of the Gated Linear Unit (GLU) (Dauphin et al., 2017). The final representation is defined as a Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$. A visualization of the proposed architecture can be found in Fig. 5.
@@
Our key hypothesis is that the Hadamard representation can prevent saturation, hence alleviating vanishing gradients and dying neurons (See Eq. 1). To support our hypothesis, we investigate the derivative of a product of two functions. For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $\uline{\color[rgb]{0.40625,0.1953125,0.97265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40625,0.1953125,0.97265625}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g^{\prime}(x)h(x)}}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g(x)h^{\prime}(x)}}$. In the context of using a sigmoid activation function for $f(x)$, the derivative of $z(x)$ becomes:
$z^{\prime}(x)=\uline{\color[rgb]{0.40625,0.1953125,0.97265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40625,0.1953125,0.97265625}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
$+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))}}$
@@
$z^{\prime}(x)$ $=\uline{\color[rgb]{0.40625,0.1953125,0.97265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40625,0.1953125,0.97265625}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{1}\,\text{sech}^{2}(A_{1}x+B_{1})\,\tanh(A_{2}x+B_{2})}}$
$\quad+\,\uline{\color[rgb]{0.296875,0.76171875,0.83203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{2}\,\text{sech}^{2}(A_{2}x+B_{2})\,\tanh(A_{1}x+B_{1})}}.$
In this context, $\operatorname{sech}^{2}$ is the derivative of the hyperbolic tangent function. Unlike the sigmoid, the hyperbolic tangent saturates to nonzero values, ensuring that if and only if both parts are saturated, product saturation occurs. Thus, when $g(x)$ saturates, $h(x)$ still keeps a non-trivial gradient in the product, providing a mechanism to avoid vanishing gradients. We visualize the kernel densities during training with a Hadamard representation in Fig. 6. The individual representations before taking the Hadamard product can be found in Appendix B.
@@
Hyperbolic Tangent:
In the case of the hyperbolic tangent, product saturation only occurs if strictly both neurons are saturated. This results in a probability of $\mathbf{p}\cdot\mathbf{p}=\mathbf{p}^{2}$. Taking a product of hyperbolic tangent activated neurons thus reduces the neuron saturation probability from $\mathbf{p}$ to $\mathbf{p}^{2}$.
Sigmoid:
Rectified Linear-Unit:
In the case of a ReLU activation, the probability of a single neuron dying is taken as $\mathbf{p}$. As we look at the product of two neurons, the probability that one of the two neurons does not saturate is therefore $1-\mathbf{p}$, and the probability that both neurons do not saturate is $(1-\mathbf{p})^{2}$. The probability that at least one neuron saturates is thus equal to $1-(1-\mathbf{p})^{2}=2\mathbf{p}-\mathbf{p}^{2}$. As the ReLU saturation results in strict zeroes, this results in the product also being zero. Taking a product of ReLU activated neurons therefore increases the final neuron saturation probability from $\mathbf{p}$ to $2\mathbf{p}-\mathbf{p}^{2}$. For an overview, we refer the reader to both Table 1 and the corresponding empirical evidence in Appendix C.3.
Activation Prob. Prob. with HR $\Delta$
Tanh $\mathbf{p}$ $\mathbf{p}^{2}$ $-(\mathbf{p}-\mathbf{p}^{2})$
Sigmoid $\mathbf{p}$ $\mathbf{p}$ 0
ReLU $\mathbf{p}$ $2\mathbf{p}-\mathbf{p}^{2}$ $+(\mathbf{p}-\mathbf{p}^{2})$
5 Experiments